OpenStudy (anonymous):

am i right? Electricity is consumed between the hours of midnight (t = 0) and 6 a.m. (t = 6). Selected values of the rate of consumption are shown in the table below, with t measured in hours past midnight, and R(t) measured in KW per hour. Use 4 trapezoids to estimate the total amount of electricity to the nearest KW consumed between midnight and 6 a.m. t (hours) 0 1 2 5 6 R(t) KW/hr 400 200 350 200 150

OpenStudy (anonymous):

@ganeshie8 @welshfella @nincompoop

OpenStudy (anonymous):

im getting 1575

OpenStudy (anonymous):

Left point sum (400 + 200 + 1050 + 200) = 1850 Right point sum (200 + 350 + 600 + 150) = 1300 Average sum (1850 - 1300)/2 = 1575

OpenStudy (anonymous):

@aaronq @pooja195 @Michele_Laino @paki @Zale101 @goformit100 does this look right to you gus?

OpenStudy (anonymous):

guys*